What is the pH of a .62 M [C5H5NH]NO3 solution?

Kb(C5H5N) = 1.7 x 10-9

Kb(NH3) = 1.8 x 10-5
Answer #1

Dear Student

Given: Molar of pyridinium nitrate [CG H g NH] NOz solution = 0.62m Ko (CS HsN) = 1,78109 Ko CNHz) = 1.8x100s Now [CS H s NH]i 3) XP 0.62 x 5.882 8106 = 3.6468x10-6 2 = 1-9097 X10-3 [H30+] = 1,9097x10-3 pH = -log [Hot] E-log (1.4092x10-3) pH = 2.719Please give a positive rating to this answer. Thank you!